OpenStudy (anonymous):

How many liters of 80% alcohol solution and 20% alcohol solution must be mixed to obtain 9 liters of 40% alcohol solution?

OpenStudy (dumbcow):

 the table gives 2 equations: x + y = 9 80x + 20y = 40(9) = 360